An account bearing interest at 6 percent compounded quarterly was established 15 years ago. The account balance now is $ 10,000. What was the initial amount when the account was established ?
1
Expert's answer
2021-08-25T17:25:28-0400

"A=P(1+\frac{r}{n})^{nt}"


"10,000=P(1+\frac{0.06}{4})^{4\times15}"


"10,000=P(1+0.015)^{60}"


"10,000=2.443219776P"


"\\P=\frac{10,000}{2.443219776}"


"P=4,092.959667"


Therefore the initial amount was $4,092.959667